# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 2.2. Coupling with Gaussian log-correlated fields
Let $\gamma\in(1,\sqrt{2})$, fix $(a_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$ i.i.d satisfying (1.2), and let $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ be defined as in (2.3). Construct $(\widetilde{g}_{n}^{(i)}(v))_{i\in\{1,2\},v\in\mathcal{N}_{n}}$ as in Eq. 2.24. Then,
$\widetilde{\mu}_{\gamma,\widetilde{a}}\ll\widetilde{\mu}_{\gamma,\widetilde{g}}\ll\widetilde{\mu}_{\gamma,\widetilde{a}}\qquad\text{almost surely.}$. (2.25)
The proof of Proposition 2.7 is completed in Section 3.
2.3. Proof of 1.1
In Proposition 2.3 and Proposition 2.7, we constructed on the same probability space $(a_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$, $(\widetilde{a}_{n})_{n\in\mathbb{N}}$, $(g_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$, and $(\widetilde{g}_{n}=\widetilde{g}_{n}^{(1)}+\widetilde{g}_{n}^{(2)})_{n\in\mathbb{N}}$ such that
$\widetilde{\mu}_{\gamma,\tilde{g}}\ll\mu_{\gamma,a}\ll\widetilde{\mu}_{\gamma,\tilde{g}}\qquad\text{almost surely.}$.
Proposition 2.3 applied with $a_{k}^{(i)}=g_{k}^{(i)}$ (and thus $\widetilde{a}_{n}=\widetilde{g}_{n}$) yields
$\mu_{\gamma,g}\ll\widetilde{\mu}_{\gamma,\tilde{g}}\ll\mu_{\gamma,g}\qquad\text{almost surely.}$.
The above two displays then yield 1.1 . ∎
@@
We have already constructed a coupling between $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ with Gaussians $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ in Section 2.2. Let $(\Omega,\mathcal{F},\mathbb{P})$ denote the common probability space on which these variables are constructed. Towards Proposition 2.7, it remains to show the absolute continuity (2.25). The natural candidate for a Radon-Nikodym derivative between $\widetilde{\mu}_{\gamma,\widetilde{a}}$ and $\widetilde{\mu}_{\gamma,\widetilde{g}}$ is the $n\to\infty$ limit of
@@
Our strategy is then to show that the following hold $\mathbb{P}$-almost-surely:
@@
Define $\widetilde{R}_{\infty,\gamma}(t):=\lim\limits_{n\to\infty}\widetilde{R}_{n,\gamma}(t)$, and note that by (A) this limit exists for $\widetilde{\mu}_{\gamma,\widetilde{a}}$-almost every $t\in[0,1]$. This along with (C) allows us to apply Vitali’s convergence theorem, which yields $\widetilde{R}_{\infty,\gamma}\in L^{1}(\widetilde{\mu}_{\gamma,\widetilde{a}})$ and, for each $A\in\mathcal{B}([0,1])$,
@@ 3.1. Proof of (A): convergence of the Radon-Nikodym derivative
Equation (2.9) implies $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)\widetilde{Z}_{n,\gamma,\widetilde{g}}(t)^{-1}$ converges in $(0,\infty)$ as $n\to\infty$ for every $t\in[0,1]$. Thus, we need only show the convergence of $e^{\gamma(\widetilde{S}_{n,\widetilde{g}}(t)-\widetilde{S}_{n,\widetilde{a}}(t))}$. This will come from the coupling estimate along thick points (Lemma 2.6), as well as Lemma 3.1 below, which states that the chaos measures are carried by thick points. Towards this, set
@@
Lemma 3.1 is proved in Appendix A. We say $t\in\mathcal{T}_{n,\gamma}$ if and only if $v_{n}(t)\in\mathcal{T}_{n,\gamma}$. Recall also the notation set forth in the display above (2.6). By definition, we have
@@
Thus, Lemma 3.1 implies it suffices to show $\widetilde{R}_{n,\gamma}(t)$ converges for $t\in\bigcup_{N\in\mathbb{N}}\bigcap_{n\geq N}\mathcal{T}_{n,\gamma}$. Let $t\in\bigcup_{N\in\mathbb{N}}\bigcap_{n\geq N}\mathcal{T}_{n,\gamma}$ be arbitrary. There exists exists a $N_{t}(\omega)\in\mathbb{N}$ such that $t\in\bigcap_{n\geq N_{t}(\omega)}\mathcal{T}_{n,\gamma}$. Furthermore, Eq. 2.20 and Borel-Cantelli imply that for each $\omega\in\Omega$, there is an $N_{0}(\omega)$ such that for all $n\geq N_{0}(\omega)$ and $t\in\mathcal{T}_{n,\gamma}$ we have $|\widetilde{g}_{n}(t)-\widetilde{a}_{n}(t)|\leq n^{-2}$. Thus, for $n\geq\max\{N_{0},N_{t}(\omega)\}$, we have
$\sum_{m=1}^{n}|\widetilde{g}_{m}(t)-\widetilde{a}_{m}(t)|\leq\sum_{m=1}^{\max\{N_{0}(\omega),N_{t}(\omega)\}}\left|\widetilde{g}_{m}(t)-\widetilde{a}_{m}(t)\right|+\sum_{m=\max\{N_{0}(\omega),N_{t}(\omega)\}}^{n}m^{-2}\leq C_{t}(\omega)<\infty\,.$
@@ 3.2. Proof of (B): integral limits
We only show the first equation (3.2), as the proof of (3.3) can be done analogously.
@@ 3.2.1. Reduction, and overview of the proof
We begin by reducing (3.2) to more tractable limit statements, providing an overview of the proof strategy along the way.
@@
It is enough to show (3.2) for sets the form $A:=I_{m}(v)$, for all $m\in\mathbb{N}$ and $v\in\mathcal{N}_{m}$. Thus, we fix $m\in\mathbb{N}$ and $v\in\mathcal{N}_{m}$, and set $\mathcal{N}_{n}(v)$ to be the set of particles in generation $n$ who are descendants of $v$. Using the fact that $\widetilde{R}_{n,\gamma}(t)$ is constant on each of the $I_{n}(w)$, $w\in\mathcal{N}_{n}(v)$, we get
@@
$\lim\limits_{n\to\infty}\int_{I_{m}(v)}\widetilde{R}_{n,\gamma}(t)\mathrm{d}\widetilde{\mu}_{\gamma,\widetilde{a}}(t)$ $=\lim\limits_{n\to\infty}\sum_{w\in\mathcal{N}_{n}(v)}\Big{(}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(w)]e^{-\gamma\widetilde{S}_{n,\widetilde{a}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{a}}(w)\Big{)}e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\,.$
@@
Since $|I_{n}(w)|\approx|\mathcal{N}_{n}(v)|$ and $\mathbb{E}[(M_{n,\widetilde{a}}(w)-1)e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}]=0$ due to the aforementioned independence, the above roughly corresponds to a law of large numbers. However, all terms in the above sum are correlated to various degrees. We prove Eq. 3.4 by moment bounds on the sum that give the desired convergence upon applying Markov’s inequality and Borel-Cantelli. An issue with this approach is that the first moment does not decay, while $\mathbb{E}[M_{n,\widetilde{a}}(w)^{2}]=\infty$ and $\mathbb{E}[(e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1})^{2}]$ grows too fast with $n$ (in light of Eq. 2.9). To deal with this, we truncate both objects by instituting “barrier events”. We need a barrier both for $\widetilde{S}_{s,\widetilde{g}}(w)$, $s\leq n$, and for $\widetilde{S}_{(n,n+r],\widetilde{a}}(t)$, which is the field below generation $n$. The differences between the truncated and un-truncated quantities can be shown to be small in $L^{1}$ by mimicking the computations done in [BerestyckiSimplePath].
@@
where the $\mathbb{P}$-almost sure limit in (3.5) exists, since
@@
is a positive supermartingale, and the $\mathbb{P}$-a.s. limit in (3.6) exists, since it is the limit of the difference between the positive martingale $(|I_{n}(w)|^{-1}\int_{I_{n}(w)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t)_{r\in\mathbb{N}_{0}}$ and the positive supermartingale in (3.7). We note that, by definition, $M_{n,\widetilde{a}}^{(1)}(w)+M_{n,\widetilde{a}}^{(2)}(w)=M_{n,\widetilde{a}}(w)$. For $i\in\{1,2\}$, we define the centered variants of these:
@@
$B_{h,n}^{\alpha}(w)$ :=$:=\big{\{}\forall s\in\{\log(n)^{1.5}+1,\dots,n\}\,:\,\widetilde{S}_{s,\widetilde{g}}(w)\leq\alpha\log(2)s\big{\}}\,,\text{ and}$
$A_{h}^{\alpha}(w)$ :=$:=\{\widetilde{S}_{h,\widetilde{g}}(w)>\alpha\log(2)h\}\,.$
Finally, recalling that $\mathbb{E}[M_{n,a}^{(1)}(w)]+\mathbb{E}[M_{n,a}^{(2)}(w)]=\mathbb{E}[M_{n,a}(w)]=1$ for all $w\in\mathcal{N}_{n}(v)$, Equation (3.4) (and thus Eq. 3.2) follows from the existence of some $\alpha\in(\gamma,\sqrt{2})$ such that:
@@
Eq. 3.8 considers a barrier on the increment of the field $\widetilde{S}_{\cdot,\widetilde{g}}(w)$ both before and after time $n$. It will be shown by controlling the second moment of the sum, and its proof will occupy the bulk of the remainder of the article. Eqs. 3.9 and 3.10 consider the increment of the field crossing a barrier before time $n$ and after time $n$ respectively; each will be shown by controlling the first moment.
3.2.2. Proofs of Eqs. 3.8, 3.9 and 3.10
@@
Before proving the lemmata, we show how they imply (3.8)–(3.10) (in reverse order). We repeatedly use the independence of $(M^{(i)}_{n,\widetilde{a}}(w))_{i\in\{1,2\},w\in\mathcal{N}_{n}(v)}$ from $(e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1})_{w\in\mathcal{N}_{n}(v)}$, referring to this fact as “generational independence”.
Proof of Eq. 3.10.
Using $\mathbb{E}[e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}]=1$, generational independence, and Lemma 3.4, we obtain the first moment bound
@@
Since $m$ is fixed, the Markov inequality and Borel-Cantelli yield Eq. 3.10. ∎
Proof of Eq. 3.9.
@@
$2|I_{m}(v)|\max_{w\in\mathcal{N}_{n}(v)}\sum_{l=\log(n)^{1.5}+1}^{n}\mathbb{E}\Big{[}\mathbbm{1}_{A_{l}^{\alpha}(w)}e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\Big{]}\\ =2|I_{m}(v)|\max_{w\in\mathcal{N}_{n}(v)}\sum_{l=\log(n)^{1.5}+1}^{n}\mathbb{E}\Big{[}\mathbbm{1}_{A_{l}^{\alpha}(w)}e^{\gamma\widetilde{S}_{l,\widetilde{g}}(w)}\widetilde{Z}_{l,\gamma,\widetilde{g}}(w)^{-1}\Big{]}\,,$ (3.12)
where the equality comes from the independence of $\widetilde{S}_{n,\widetilde{g}}(w)-\widetilde{S}_{l,\widetilde{g}}(w)$ from $\widetilde{S}_{l,\widetilde{g}}(w)$, which also gives
$\mathbb{E}\Big{[}e^{\gamma(\widetilde{S}_{n,\widetilde{g}}(w)-\widetilde{S}_{l,\widetilde{g}}(w))}\Big{]}=\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)\widetilde{Z}_{l,\gamma,\widetilde{g}}(w)^{-1}\,.$
@@
$\mathbbm{1}_{A_{l}^{\alpha}(w)}\leq e^{(\alpha-\gamma)\widetilde{S}_{l,\widetilde{g}}(w)-(\alpha-\gamma)\alpha\log(2)l}.$ (3.13)
Plugging Lemma 2.2 and Eq. 3.13 into Eq. 3.12 yields
$\mathbb{E}\bigg{|}\sum_{w\in\mathcal{N}_{n}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\bigg{|}\\ \begin{aligned} &\leq C\max_{w\in\mathcal{N}_{n}(v)}\sum_{l=\log(n)^{1.5}+1}^{n}\mathbb{E}\big{[}e^{\alpha\widetilde{S}_{l,\widetilde{g}}(w)}\big{]}e^{-(\alpha-\gamma)\alpha\log(2)l-\frac{\gamma^{2}}{2}\log(2)l}\\ &\leq C\sum_{l=\log(n)^{1.5}+1}^{n}2^{-\frac{(\alpha-\gamma)^{2}}{2}l}\leq Ce^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{1.5}}\,.\end{aligned}$
Eq. 3.9 now follows from the Markov inequality and Borel-Cantelli. ∎
Proof of Eq. 3.8.
@@
$\widetilde{E}_{n,\gamma,\widetilde{g}}:=e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\,,\qquad\widetilde{E}_{(k,n],\gamma,\widetilde{g}}:=e^{\gamma\widetilde{S}_{(k,n],\widetilde{g}}(w)}\widetilde{Z}_{(k,n],\gamma,\widetilde{g}}(w)^{-1}.$
We proceed by bounding the second moment of the sum in Eq. 3.8. Our strategy will be to repartition the sum over $\mathcal{N}_{n}(v)$ such that the $\overline{M}_{n,\widetilde{a}}^{(1)}(w)$ in each partition have small correlation. Towards this, keeping in mind Lemma 3.5, we write the disjoint union

### v2

@@ 2.2. Coupling with Gaussian log-correlated fields
Let $\gamma\in(1,\sqrt{2})$, fix $(a_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$ i.i.d satisfying Eq.˜1.2, and let $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ be defined as in (2.3). Construct $(\widetilde{g}_{n}^{(i)}(v))_{i\in\{1,2\},v\in\mathcal{N}_{n}}$ as in Eq.˜2.24. Then,
$\widetilde{\mu}_{\gamma,\widetilde{a}}\ll\widetilde{\mu}_{\gamma,\widetilde{g}}\ll\widetilde{\mu}_{\gamma,\widetilde{a}}\qquad\text{almost surely.}$ (2.25)
The proof of Proposition˜2.7 is completed in Section˜3.
2.3. Proof of ˜1.1
In Proposition˜2.3 and Proposition˜2.7, we constructed on the same probability space $(a_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$, $(\widetilde{a}_{n})_{n\in\mathbb{N}}$, $(g_{k}^{(i)})_{i\in\{1,2\},k\in\mathbb{N}}$, and $(\widetilde{g}_{n}=\widetilde{g}_{n}^{(1)}+\widetilde{g}_{n}^{(2)})_{n\in\mathbb{N}}$ such that
$\widetilde{\mu}_{\gamma,\tilde{g}}\ll\mu_{\gamma,a}\ll\widetilde{\mu}_{\gamma,\tilde{g}}\qquad\text{almost surely.}$
Proposition˜2.3 applied with $a_{k}^{(i)}=g_{k}^{(i)}$ (and thus $\widetilde{a}_{n}=\widetilde{g}_{n}$) yields
$\mu_{\gamma,g}\ll\widetilde{\mu}_{\gamma,\tilde{g}}\ll\mu_{\gamma,g}\qquad\text{almost surely.}$
The above two displays then yield ˜1.1 . ∎
@@
We have already constructed a coupling between the non-Gaussian random variables $(\widetilde{a}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ with Gaussians $(\widetilde{g}_{n}^{(i)})_{i\in\{1,2\},n\in\mathbb{N}}$ in Section˜2.2. Let $(\Omega,\mathcal{F},\mathbb{P})$ denote the common probability space on which these variables are constructed. Towards Proposition˜2.7, it remains to show the absolute continuity Eq.˜2.25. The natural candidate for a Radon-Nikodym derivative between $\widetilde{\mu}_{\gamma,\widetilde{a}}$ and $\widetilde{\mu}_{\gamma,\widetilde{g}}$ is the $n\to\infty$ limit of
@@
Our strategy is then to show that the following hold $\mathbb{P}$-almost surely:
@@
[…]
@@ 3.1. Proof of (A): convergence of the Radon-Nikodym derivative
Equation˜2.9 implies $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)\widetilde{Z}_{n,\gamma,\widetilde{g}}(t)^{-1}$ converges in $(0,\infty)$ as $n\to\infty$ for every $t\in[0,1]$. Thus, we need only show the convergence of $e^{\gamma(\widetilde{S}_{n,\widetilde{g}}(t)-\widetilde{S}_{n,\widetilde{a}}(t))}$. This will come from the coupling estimate along thick points (Lemma˜2.6), as well as Lemma˜3.1 below, which states that the chaos measures are carried by thick points. Towards this, set
@@
Lemma˜3.1 is proved in Appendix A. We say $t\in\mathcal{T}_{n,\gamma}$ if and only if $v_{n}(t)\in\mathcal{T}_{n,\gamma}$. Recall also the notation set forth in the display above (2.6). By definition, we have
@@
Thus, Lemma 3.1 implies it suffices to show $\widetilde{R}_{n,\gamma}(t)$ converges for $t\in\bigcup_{N\in\mathbb{N}}\bigcap_{n\geq N}\mathcal{T}_{n,\gamma}$. Let $t\in\bigcup_{N\in\mathbb{N}}\bigcap_{n\geq N}\mathcal{T}_{n,\gamma}$ be arbitrary. There exists exists a $N_{t}(\omega)\in\mathbb{N}$ such that $t\in\bigcap_{n\geq N_{t}(\omega)}\mathcal{T}_{n,\gamma}$. Furthermore, Eq.˜2.20 and Borel-Cantelli imply that for each $\omega\in\Omega$, there is an $N_{0}(\omega)$ such that for all $n\geq N_{0}(\omega)$ and $t\in\mathcal{T}_{n,\gamma}$ we have $|\widetilde{g}_{n}(t)-\widetilde{a}_{n}(t)|\leq n^{-2}$. Set $\widetilde{N}_{t}(\omega):=\max\{N_{0}(\omega),N_{t}(\omega)\}$. For $n\geq\widetilde{N}_{t}(\omega)$, we have
$\sum_{m=1}^{n}|\widetilde{g}_{m}(t)-\widetilde{a}_{m}(t)|\leq\sum_{m=1}^{\widetilde{N}_{t}(\omega)}\left|\widetilde{g}_{m}(t)-\widetilde{a}_{m}(t)\right|+\sum_{m=\widetilde{N}_{t}(\omega)}^{n}m^{-2}\leq C_{t}(\omega)<\infty\,.$
@@ 3.2. Proof of (B): integral limits
We only show the first equation Eq.˜3.2, as the proof of Eq.˜3.3 can be done analogously.
@@ 3.2.1. Reduction, and overview of the proof
We begin by reducing Eq.˜3.2 to more tractable limit statements, providing an overview of the proof strategy along the way.
@@
It suffices to show Eq.˜3.2 for sets of the form $A:=I_{m}(v)$, for all $m\in\mathbb{N}$ and $v\in\mathcal{N}_{m}$. Thus, we fix $m\in\mathbb{N}$ and $v\in\mathcal{N}_{m}$, and set $\mathcal{N}_{n}(v)$ to be the set of particles in generation $n$ who are descendants of $v$. Using that $\widetilde{R}_{n,\gamma}(t)$ is constant on each of the $I_{n}(w)$, $w\in\mathcal{N}_{n}(v)$, we get
@@
$\lim\limits_{n\to\infty}\int_{I_{m}(v)}\widetilde{R}_{n,\gamma}(t)\mathrm{d}\widetilde{\mu}_{\gamma,\widetilde{a}}(t)$ $=\lim\limits_{n\to\infty}\sum_{w\in\mathcal{N}_{n}(v)}\Big(\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(w)]e^{-\gamma\widetilde{S}_{n,\widetilde{a}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{a}}(w)\Big)$
$\quad\qquad\times e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\,.$
@@
Since $|I_{n}(w)|\approx|\mathcal{N}_{n}(v)|^{-1}$ and due to the aforementioned independence $\mathbb{E}[(M_{n,\widetilde{a}}(w)-1)e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}]=0$, the above roughly corresponds to a law of large numbers. However, all terms in the above sum are correlated to various degrees. We prove Eq.˜3.4 by moment bounds on the sum that give the desired convergence upon applying Markov’s inequality and Borel-Cantelli. An issue with this approach is that the first moment does not decay, while $\mathbb{E}[M_{n,\widetilde{a}}(w)^{2}]=\infty$ and $\mathbb{E}[(e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1})^{2}]$ grows too fast with $n$ (in light of Eq.˜2.9). To deal with this, we truncate both objects by instituting “barrier events”. We need a barrier both for $\widetilde{S}_{s,\widetilde{g}}(w)$, $s\leq n$, and for $\widetilde{S}_{(n,n+r],\widetilde{a}}(t)$, which is the field below generation $n$. The differences between the truncated and un-truncated quantities can be shown to be small in $L^{1}$ by mimicking the computations done in [6].
@@
where the $\mathbb{P}$-almost sure limit in Eq.˜3.5 exists, since
@@
is a positive supermartingale. The $\mathbb{P}$-a.s. limit in Eq.˜3.6 exists, since it is the limit of the difference between the positive martingale $(|I_{n}(w)|^{-1}\int_{I_{n}(w)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}/\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)\;\mathrm{d}t)_{r}$ and the positive supermartingale in Eq.˜3.7. We note that, by definition, $M_{n,\widetilde{a}}^{(1)}(w)+M_{n,\widetilde{a}}^{(2)}(w)=M_{n,\widetilde{a}}(w)$. For $i\in\{1,2\}$, we define the centered variants of these:
@@
$B_{h,n}^{\alpha}(w)$ $:=\big\{\forall s\in\{h+1,\dots,n\}\,:\,\widetilde{S}_{s,\widetilde{g}}(w)\leq\alpha\log(2)s\big\}\,,\text{ and}$
$A_{h}^{\alpha}(w)$ $:=\{\widetilde{S}_{h,\widetilde{g}}(w)>\alpha\log(2)h\}\,.$
Finally, recalling that $\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w)]+\mathbb{E}[M_{n,\widetilde{a}}^{(2)}(w)]=\mathbb{E}[M_{n,\widetilde{a}}(w)]=1$ for all $w\in\mathcal{N}_{n}(v)$, Eq.˜3.4 (and thus Eq.˜3.2) follows from the existence of some $\alpha\in(\gamma,\sqrt{2})$ such that $\mathbb{P}$-a.s.:
@@
Equation˜3.8 considers a barrier on the increments of the field $\widetilde{S}_{\cdot,\widetilde{g}}(w)$ before level $n$ from the $B^{\alpha}_{\log(n)^{1.5},n}(w)$ term, as well as a barrier on the increments of $\widetilde{S}_{(n,\,\cdot\,],\widetilde{a}}(w)$ after level $n$ from the $\overline{M}_{n,\widetilde{a}}^{(1)}(w)$ term. It will be shown by controlling the second moment of the sum, and its proof will occupy the bulk of the remainder of the article. Equations˜3.9 and 3.10 consider the increment of the field $\widetilde{S}_{\cdot,\widetilde{g}}(w)$ crossing a barrier before level $n$ respectively the increment of the field $\widetilde{S}_{(n,\,\cdot\,],\widetilde{a}}(w)$ crossing a barrier after level $n$; each will be shown by controlling the first moment.
3.2.2. Proofs of Eqs.˜3.8, 3.9 and 3.10
@@
Before proving the lemmata, we show how they imply Eqs.˜3.8, 3.9 and 3.10 (in reverse order). We use the independence of $(M^{(i)}_{n,\widetilde{a}}(w))_{i\in\{1,2\},w\in\mathcal{N}_{n}(v)}$ from $(e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1})_{w\in\mathcal{N}_{n}(v)}$ repeatedly, referring to this fact as “generational independence”.
Proof of Eq.˜3.10.
Using $\mathbb{E}[e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}]=1$, generational independence, as well as Lemma 3.4, we obtain the first moment bound
@@
Since $m$ is fixed, the Markov inequality and Borel-Cantelli yield Eq.˜3.10. ∎
Proof of Eq.˜3.9.
@@
$2|I_{m}(v)|\max_{w\in\mathcal{N}_{n}(v)}\sum_{\ell=\log(n)^{1.5}+1}^{n}\mathbb{E}\Big[\mathbbm{1}_{A_{\ell}^{\alpha}(w)}e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\Big]\\ =2|I_{m}(v)|\max_{w\in\mathcal{N}_{n}(v)}\sum_{\ell=\log(n)^{1.5}+1}^{n}\mathbb{E}\Big[\mathbbm{1}_{A_{\ell}^{\alpha}(w)}e^{\gamma\widetilde{S}_{\ell,\widetilde{g}}(w)}\widetilde{Z}_{\ell,\gamma,\widetilde{g}}(w)^{-1}\Big]\,$, (3.12)
where the equality comes from the independence of $\widetilde{S}_{n,\widetilde{g}}(w)-\widetilde{S}_{\ell,\widetilde{g}}(w)$ from $\widetilde{S}_{\ell,\widetilde{g}}(w)$, which also gives
$\mathbb{E}\Big[e^{\gamma(\widetilde{S}_{n,\widetilde{g}}(w)-\widetilde{S}_{\ell,\widetilde{g}}(w))}\Big]=\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)\widetilde{Z}_{\ell,\gamma,\widetilde{g}}(w)^{-1}\,.$
@@
$\mathbbm{1}_{A_{\ell}^{\alpha}(w)}\leq e^{(\alpha-\gamma)\widetilde{S}_{\ell,\widetilde{g}}(w)-(\alpha-\gamma)\alpha\log(2)\ell}.$ (3.13)
Plugging Lemma˜2.2 and Eq.˜3.13 into Eq.˜3.12 yields
$\mathbb{E}\bigg|\sum_{w\in\mathcal{N}_{n}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\mathbbm{1}_{B_{\log(n)^{1.5},n}^{\alpha}(w)^{c}}\bigg|\\ \begin{aligned} &\leq C\max_{w\in\mathcal{N}_{n}(v)}\sum_{\ell=\log(n)^{1.5}+1}^{n}\mathbb{E}\big[e^{\alpha\widetilde{S}_{\ell,\widetilde{g}}(w)}\big]e^{-(\alpha-\gamma)\alpha\log(2)\ell-\frac{\gamma^{2}}{2}\log(2)\ell}\\ &\leq C\sum_{\ell=\log(n)^{1.5}+1}^{n}2^{-\frac{(\alpha-\gamma)^{2}}{2}\ell}\leq Ce^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{1.5}}\,.\end{aligned}$
Equation˜3.9 now follows from the Markov inequality and Borel-Cantelli. ∎
Proof of Eq.˜3.8.
@@
$\widetilde{E}_{n,\gamma,\widetilde{g}}(w):=e^{\gamma\widetilde{S}_{n,\widetilde{g}}(w)}\widetilde{Z}_{n,\gamma,\widetilde{g}}(w)^{-1}\,,\qquad\widetilde{E}_{(k,n],\gamma,\widetilde{g}}(w):=e^{\gamma\widetilde{S}_{(k,n],\widetilde{g}}(w)}\widetilde{Z}_{(k,n],\gamma,\widetilde{g}}(w)^{-1},$
for $\in\mathbb{N}, knk\leq n$, $w\in\mathcal{N}_{n}$. We analogously define $\widetilde{E}_{n,\gamma,\widetilde{a}}(w)$, $\widetilde{E}_{(k,n],\gamma,\widetilde{a}}(w)$.
We proceed by bounding the second moment of the sum in Eq.˜3.8. Our strategy will be to repartition the sum over $\mathcal{N}_{n}(v)$ such that the $\overline{M}_{n,\widetilde{a}}^{(1)}(w)$ in each partition have small correlation. Towards this, keeping in mind Lemma 3.5, we write the disjoint union